# Aspects of a novel nonlinear electrodynamics in flat spacetime and in a gravity-coupled scenario

### v1

@@ I Introduction
In Maxwell’s electrodynamics, the electric field and self-energy of a point charge exhibit a divergence as the radial distance approaches zero. One of the well-known procedures to address this issue involves modifying Maxwell’s linear Lagrangian in a nonlinear manner. Introducing nonlinearity in the Lagrangian has implications for the field equations, which may or may not exhibit the linear Maxwell limit. In the 1930s, Born and Infeld first developed their nonlinear model, largely known as Born-Infeld (BI) electrodynamics [1, 2]. The classical BI Lagrangian imposes a maximum limit on the intensity of the electric field, preventing it from becoming infinite at small distances. As a result, the energy of a point electric charge at its location becomes finite. Subsequently, several other exciting nonlinear modifications of Maxwell’s Lagrangian have emerged in the literature, collectively referred to as ‘nonlinear electrodynamics’ (NLE). For example, the effective Euler-Heisenberg Lagrangian [3] originating from the 1-loop QED calculation and the one-parameter family of ModMax NLE theory [4] are worth mentioning as viable alternatives. Recently, there have been various advancements in constructing novel NLE models, all of which aim to resolve the issue of singularity in the linear model [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15], with notable features. Each of them can be identified by a parameter $\eta$. When $\eta\to 0$ or $\eta\to\infty$, depending on the formulation, the original linear theory of Maxwell is restored.
@@
There is widespread anticipation that quantum extensions of classical theories would remove or resolve the singularities present in those theories [20]. For instance, the divergence problem of Maxwell’s electrodynamics can be cured by invoking quantum electrodynamics [21]. On the other hand, as discussed above, the classical NLE models may also address the problem. There are some effective models at various levels of string/M-theory too [22, 23]. Similar to linear electrodynamics, the singularity issue in classical GR may be approached in different ways. One possible direction is to invoke ‘quantum gravity’, the quantum theory of classical GR, which is still in a state of ‘progress’. Second, one can modify Einstein’s GR (by deforming the Einstein-Hilbert action) and look for a divergence-free solution in that specific theory of gravity. An alternative method involves exploring various nonlinear generalisations of Einstein-Maxwell theory. Thus, it is expected that the problem of spacetime singularities may be resolved by minimal coupling of a suitable NLE Lagrangian with gravity [24, 25]. Our work primarily revolves around this line of thought.
@@
First, in 1968, Bardeen came up with his ad hoc proposal of a singularity-free black hole solution [26]. Subsequently, several suggestions of regular models have been reported in scientific publications [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. All of these black holes have been often referred to as ‘regular black holes’. As indicated earlier, it is found that one of the ways to understand regular solutions is the minimal coupling of Einstein’s GR with a chosen version of the matter NLE Lagrangian [38, 39, 40, 41, 42, 43, 44].
@@
The derivation of matter NLE Lagrangians usually involves reverse engineering of already proposed regular models, leaving them open to criticism from various angles. First, the dynamics of the NLE Lagrangians need to be clearly understood, at least in flat spacetime. The nonlinear evolution of waves, their scattering, birefringence, and other related effects of the NLEs are still not explored. Second, most of the NLE models contain a fractional power of the Maxwell invariant ($F=\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$), which limits the consideration to only positive values of $F$. It is not an issue when constructing regular black holes because a nonlinear magnetic monopole (where $F$ is positive) can sustain them. However, the analysis of gravitational coupling of these NLEs sourced by an electric monopole ($F$ is negative) is not doable. Since electric charge ($F$ is negative) exists in nature, the NLE models are not fully physically reasonable. Last but not the least, there are examples of electrically charged regular black holes [45, 46, 47], whose matter Lagrangian violates the Maxwellian limit, and the Lagrangian description is ambiguous (multi-valued function).
@@
Considering the issues mentioned above, we are motivated to construct a novel NLE Lagrangian model which follows the Maxwellian limit and is free from any fractional power of $F$. In other words, we aim to examine its coupling with gravity in order to create black hole solutions with both electric and magnetic characteristics. Before introducing a curved spacetime background, we analyse our proposed model in flat spacetime to establish its credibility as an alternative to the BI model. There is already a significant amount of research on this subject, specifically discussing the role of NLE models in sourcing various types of spacetime solutions [48, 49, 50, 51].
@@
As discussed before, the primary objective of NLE was to eliminate the self-energy singularity in linear electrodynamics. However, we must consider to what extent we can benefit from this process and whether eliminating spacetime singularities is eventually possible. We argue that the NLE model presented in this paper could cure the singularities in Maxwell’s electrodynamics in flat spacetime (as well as in the specific Schwarzschild-gauge curved geometry discussed in Section VI.2) and also generate regular spacetimes when coupled to Einstein’s GR in the presence of a suitable source charge.
@@ II Nonlinear electrodynamics model and field equations
Eq.(10) can be expressed in tensor form as $\textbf{D}_{i}=\epsilon_{i}{}^{j}\textbf{E}_{j}$, where $\epsilon_{ij}$ represents the electric permittivity tensor given below,
@@
where we have substituted $F=-E^{2}/2$. As the radial coordinate $r\in\mathbb{R}^{+}$, the above equation indicates a restriction on the domain of $F$, i.e. $F$ is larger than $-\frac{1}{\eta}$. However, there is no such restriction on $F$ for the purely magnetic case. Therefore, in general, $F$ can vary between $-\frac{1}{\eta}$ to $\infty$. One may note that in the Maxwellian limit, the boundary of $F$ extends to (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) as $\eta\to 0$, which is expected.
@@ III Energy of a point charge
One can numerically solve the above integral for different values of $\eta$. This gives us the self-energy of a point charge, which depends on the electric charge $e$ as well as the model parameters $\gamma$ and $\eta$. In Figure 3, the variation of the new variable $\bar{\mathbf{\epsilon}}\equiv\mathbf{\epsilon}\sqrt{4\pi\gamma}/e^{3/2}$, which is the scaled version of total energy $\epsilon$, with $\eta$ is presented. It is evident that for an arbitrary value of $\eta$, the energy of a point charge is finite. It becomes ‘infinite’ when $\eta\to 0$, as expected (Maxwellian limit). Thus, the electrostatic energy of a point charge is finite in our model, like it is in the well-known BI model.
@@ IV Vacuum birefringence
$d_{i}=\frac{\partial L^{(2)}}{\partial e^{i}}=(\alpha\delta_{i}^{j}+\beta\bar{E}_{i}\bar{E}^{j})e_{j}$ (32)
$h_{i}=-\frac{\partial L^{(2)}}{\partial b_{i}}=\alpha\delta_{i}^{j}b_{j}$
@@
$\epsilon_{ij}=\alpha\delta_{ij}+\beta\bar{E}_{i}\bar{E}_{j}$ (34)
$(\mu^{-1})_{ij}=\alpha\delta_{ij}$
@@
$\Lambda=\begin{bmatrix}-k^{2}\alpha+\omega^{2}(\alpha+\beta\bar{E}^{2})&0&0\\ 0&-k^{2}\alpha+\omega^{2}\alpha&0\\ 0&0&\omega^{2}\alpha\end{bmatrix}$ (38)
@@ VI.1 The magnetic regular black hole solution
$g_{tt}\to-1\hskip 28.45274ptand\hskip 28.45274ptg_{rr}\to 1\hskip 28.45274ptas\hskip 28.45274ptr\to\infty$
@@
$g_{tt}\to-(1-c^{2}r^{2})\hskip 28.45274ptand\hskip 28.45274ptg^{rr}\to(1-c^{2}r^{2})\hskip 28.45274ptas\hskip 28.45274ptr\to 0$
@@ VI.1.2 Energy conditions
Let us now look at the energy conditions (restricted by GR) for the matter required to support the above geometry. One defines $\rho=-T_{t}^{t}$, $\tau=T_{r}^{r}$ and $p=T_{\theta}^{\theta}=T_{\phi}^{\phi}$. By using Eqs. (55), (56) and substituting the specific values of the model parameters provided in Eq. (60), it is possible to express,

### v2

@@ I Introduction
In Maxwell’s electrodynamics, the electric field and self-energy of a point charge exhibit a divergence as the radial distance approaches zero. One of the well-known procedures to address this issue involves modifying Maxwell’s linear Lagrangian in a nonlinear manner. In the 1930s, Born and Infeld first developed their nonlinear model, largely known as Born-Infeld (BI) electrodynamics [1, 2]. The classical BI Lagrangian imposes a maximum limit on the intensity of the electric field, preventing it from becoming infinite at small distances. As a result, the energy of a point electric charge at its location becomes finite. Subsequently, several other exciting nonlinear modifications of Maxwell’s Lagrangian have emerged in the literature, collectively referred to as ‘nonlinear electrodynamics’ (NLE). For example, the effective Euler-Heisenberg Lagrangian [3] originating from the 1-loop QED calculation and the one-parameter family of ModMax NLE theory [4] are worth mentioning as viable alternatives. Recently, there have been various advancements in constructing novel NLE models, all of which aim to resolve the issue of singularity in the linear model [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15], with notable features. Each of them can be identified by a parameter $\eta$. When $\eta\to 0$ or $\eta\to\infty$, depending on the formulation, the original linear theory of Maxwell is restored.
@@
There is widespread anticipation that quantum extensions of classical theories would remove or resolve the singularities present in those theories [20]. For instance, the divergence problem of Maxwell’s electrodynamics can be cured by invoking quantum electrodynamics [21]. There are some effective models at various levels of string/M-theory too [22, 23]. Similar to linear electrodynamics, the singularity issue in classical GR may be approached in different ways. One possible direction is to invoke ‘quantum gravity’, the quantum theory of classical GR, which is still in a state of ‘progress’. Second, one can modify Einstein’s GR (by deforming the Einstein-Hilbert action) and look for a divergence-free solution in that specific theory of gravity. An alternative method involves exploring various nonlinear generalisations of Einstein-Maxwell theory. Thus, it is expected that the problem of spacetime singularities may be resolved by minimal coupling of a suitable NLE Lagrangian with gravity [24, 25]. Our work primarily revolves around this line of thought.
@@
First, in 1968, Bardeen came up with his ad hoc proposal of a singularity-free black hole solution [26]. Subsequently, several suggestions of regular models have been reported in scientific publications [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. All of these black holes have been often referred to as ‘regular black holes’. As indicated earlier, it is found that one of the ways to understand regular solutions is the minimal coupling of Einstein’s GR with a chosen version of the matter NLE Lagrangian [38, 39, 40, 41, 42, 43, 44]. In recent times, there have been several new constructions of regular black holes with unique features [45, 46, 47, 48, 49, 50, 51, 52, 53].
@@
The derivation of matter NLE Lagrangians usually involves reverse engineering of already proposed regular models as is done for a generic class of regular black holes in [54, 55], leaving them open to criticism from various angles. First, the dynamics of the NLE Lagrangians need to be clearly understood, at least in flat spacetime. The nonlinear evolution of waves, their scattering, birefringence, and other related effects are still not explored. Second, the matter NLE models contain a fractional power of the Maxwell invariant ($F=\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$), which limits the consideration to only positive values of $F$. As a result, the electric field of a point charge ($F$ negative) cannot be described in the context of these Lagrangians. However, it is not an issue in constructing a regular black hole because a nonlinear magnetic monopole (where $F$ is positive) can sustain it. But, the minimal coupling of these Lagrangians with GR in the presence of an electric point charge remains out of scope. As electric charge exists in nature, such NLE models seem to have limitations. Moreover, the original motivation of nonlinear electrodynamics (elimination of the singularity in the Maxwell field – as mentioned earlier) is not fulfilled by these NLE Lagrangians. Lastly, we also note that there are examples of electrically charged regular black holes [56, 57, 58], though the matter Lagrangian violates the Maxwellian limit, and the Lagrangian description is somewhat ambiguous (multi-valued functions).
@@
Considering the issues mentioned above, we are motivated to construct a novel NLE Lagrangian model which follows the Maxwellian limit and is free from any fractional power of $F$. In other words, we aim to examine its coupling with gravity in order to create black hole solutions with both electric and magnetic characteristics. Before introducing a curved spacetime background, we analyse our proposed model in flat spacetime to establish its credibility as an alternative to the BI model. There is already a significant amount of research on this subject, specifically discussing the role of NLE models in sourcing various types of spacetime solutions [59, 60, 61, 62, 63, 64, 65].
@@
As discussed before, the primary objective of NLE was to eliminate the self-energy singularity in linear electrodynamics. However, we must consider to what extent we can benefit from this process and whether eliminating spacetime singularities is eventually possible. We argue that the NLE model presented in this paper could cure the singularities in Maxwell’s electrodynamics in flat spacetime and also generate regular spacetimes when coupled to Einstein’s GR in the presence of a suitable source charge.
@@ II Nonlinear electrodynamics model and field equations
where E is the electric field. Eq.(10) can be expressed in tensor form as $\textbf{D}_{i}=\epsilon_{i}{}^{j}\textbf{E}_{j}$, where $\epsilon_{ij}$ represents the electric permittivity tensor given below,
@@
where we have substituted $F=-E^{2}/2$ and $E$ is the magnitude of electric field E. As the radial coordinate $r\in\mathbb{R}^{+}$, the above equation indicates a restriction on the domain of $F$, i.e. $F$ is larger than $-\frac{1}{\eta}$. However, there is no such restriction on $F$ for the purely magnetic case. Therefore, in general, $F$ can vary between $-\frac{1}{\eta}$ to $\infty$. One may note that in the Maxwellian limit, the boundary of $F$ extends to (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) as $\eta\to 0$, which is expected.
@@ III Energy of a point charge
One can numerically solve the above integral for different values of $\eta$. This gives us the self-energy of a point charge, which depends on the electric charge $e$ as well as the model parameters $\gamma$ and $\eta$. In Figure 3, the variation of the new variable $\bar{\mathbf{\epsilon}}\equiv\mathbf{\epsilon}\sqrt{4\pi\gamma}/e^{3/2}$, which is the scaled version of total energy $\epsilon$, with $\eta$ is presented. It is evident that for an arbitrary value of $\eta$, the energy of a point charge is finite. And in the Maxwell limit ($\eta\to 0$), the integral in Eq.(29) becomes,
$\mathbf{\epsilon}\approx\frac{e^{3/2}}{\sqrt{4\pi\gamma}}\left[\frac{\sqrt{E}}{2}\right]_{0}^{\infty}$ (30)
which shows an infinite amount of energy (as expected). Thus, the electrostatic energy of a point charge is finite in our model ($\eta\neq 0$), like it is in the well-known BI model.
@@ IV Vacuum birefringence
$d_{i}=\frac{\partial L^{(2)}}{\partial e^{i}}=(\alpha\delta_{i}^{j}+\beta\bar{E}_{i}\bar{E}^{j})e_{j},\hskip 14.22636pth_{i}=-\frac{\partial L^{(2)}}{\partial b_{i}}=\alpha\delta_{i}^{j}b_{j}$ (33)
@@
$\epsilon_{ij}=\alpha\delta_{ij}+\beta\bar{E}_{i}\bar{E}_{j},\hskip 14.22636pt(\mu^{-1})_{ij}=\alpha\delta_{ij}$ (35)
@@
$\Lambda=\begin{bmatrix}-k^{2}\alpha+\omega^{2}(\alpha+\beta\bar{E}^{2})&0&0\\ 0&-k^{2}\alpha+\omega^{2}\alpha&0\\ 0&0&\omega^{2}\alpha\end{bmatrix}$
@@ VI.1 The magnetic regular black hole solution
[…]
@@
$g_{tt}\to-(1-c^{2}r^{2})\hskip 28.45274pt\text{and}\hskip 28.45274ptg^{rr}\to(1-c^{2}r^{2})\hskip 28.45274pt\text{as}\hskip 28.45274ptr\to 0$
@@ VI.1.2 Energy conditions
Let us now look at the energy conditions (restricted by GR) for the matter required to support the above geometry. One defines $\rho=-T_{t}^{t}$, $\tau=T_{r}^{r}$ and $p=T_{\theta}^{\theta}=T_{\phi}^{\phi}$. By using Eqs. (54), (55) and substituting the specific values of the model parameters provided in Eq. (VI.1), it is possible to express,